OpenStudy (anonymous):

Complete the standard form of the equation –3y = 8 – 4x. _______x + _______y = 8

OpenStudy (anonymous):

@chosenmatt any good with math? There's like 5 questions of crap I dont know.

OpenStudy (chosenmatt):

one sec

OpenStudy (anonymous):

Alrightt

OpenStudy (chosenmatt):

4x+-3y=8

OpenStudy (anonymous):

Simplifying -4x + -3y = 6 Solving -4x + -3y = 6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3y' to each side of the equation. -4x + -3y + 3y = 6 + 3y Combine like terms: -3y + 3y = 0 -4x + 0 = 6 + 3y -4x = 6 + 3y Divide each side by '-4'. x = -1.5 + -0.75y Simplifying x = -1.5 + -0.75y MEDAL IF RIGHT

OpenStudy (anonymous):

So where'd the =6 come from. And thanks. @chosenmatt And @The_Jokers_wife

OpenStudy (chosenmatt):

well i think it would be positive 4 not negative 4

OpenStudy (chosenmatt):

but that's what i THINK xD not a fact

OpenStudy (anonymous):

@chosenmatt I used your answer lmao. We'll find out here inna minute.

OpenStudy (chosenmatt):

(nervous) XD

OpenStudy (anonymous):

Lmao hold on.

OpenStudy (anonymous):

@chosenmatt You were right.

OpenStudy (chosenmatt):

(dances in excitement)

OpenStudy (chosenmatt):

lol

OpenStudy (anonymous):

standard form meaning 4x + (-3y) = 8?

OpenStudy (anonymous):

Lol @ciara9271 sorry I cont help